### v1

# Observation of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$

Observation of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$
M. Ablikim1, M. N. Achasov4,c, P. Adlarson76, O. Afedulidis3, X. C. Ai81, R. Aliberti35, A. Amoroso75A,75C, Q. An72,58,a, Y. Bai57, O. Bakina36, I. Balossino29A, Y. Ban46,h, H.-R. Bao64, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi75A,75C, E. Bianco75A,75C, A. Bortone75A,75C, I. Boyko36, R. A. Briere5, A. Brueggemann69, H. Cai77, X. Cai1,58, A. Calcaterra28A, G. F. Cao1,64, N. Cao1,64, S. A. Cetin62A, X. Y. Chai46,h, J. F. Chang1,58, G. R. Che43, Y. Z. Che1,58,64, G. Chelkov36,b, C. Chen43, C. H. Chen9, Chao Chen55, G. Chen1, H. S. Chen1,64, H. Y. Chen20, M. L. Chen1,58,64, S. J. Chen42, S. L. Chen45, S. M. Chen61, T. Chen1,64, X. R. Chen31,64, X. T. Chen1,64, Y. B. Chen1,58, Y. Q. Chen34, Z. J. Chen25,i, Z. Y. Chen1,64, S. K. Choi10, G. Cibinetto29A, F. Cossio75C, J. J. Cui50, H. L. Dai1,58, J. P. Dai79, A. Dbeyssi18, R.  E. de Boer3, D. Dedovich36, C. Q. Deng73, Z. Y. Deng1, A. Denig35, I. Denysenko36, M. Destefanis75A,75C, F. De Mori75A,75C, B. Ding67,1, X. X. Ding46,h, Y. Ding40, Y. Ding34, J. Dong1,58, L. Y. Dong1,64, M. Y. Dong1,58,64, X. Dong77, M. C. Du1, S. X. Du81, Y. Y. Duan55, Z. H. Duan42, P. Egorov36,b, Y. H. Fan45, J. Fang1,58, J. Fang59, S. S. Fang1,64, W. X. Fang1, Y. Fang1, Y. Q. Fang1,58, R. Farinelli29A, L. Fava75B,75C, F. Feldbauer3, G. Felici28A, C. Q. Feng72,58, J. H. Feng59, Y. T. Feng72,58, M. Fritsch3, C. D. Fu1, J. L. Fu64, Y. W. Fu1,64, H. Gao64, X. B. Gao41, Y. N. Gao46,h, Yang Gao72,58, S. Garbolino75C, I. Garzia29A,29B, L. Ge81, P. T. Ge19, Z. W. Ge42, C. Geng59, E. M. Gersabeck68, A. Gilman70, K. Goetzen13, L. Gong40, W. X. Gong1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco75A,75C, M. H. Gu1,58, Y. T. Gu15, C. Y. Guan1,64, A. Q. Guo31,64, L. B. Guo41, M. J. Guo50, R. P. Guo49, Y. P. Guo12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han64, T. T. Han1, F. Hanisch3, X. Q. Hao19, F. A. Harris66, K. K. He55, K. L. He1,64, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng1,58,64, C. Herold60, T. Holtmann3, P. C. Hong34, G. Y. Hou1,64, X. T. Hou1,64, Y. R. Hou64, Z. L. Hou1, B. Y. Hu59, H. M. Hu1,64, J. F. Hu56,j, S. L. Hu12,g, T. Hu1,58,64, Y. Hu1, G. S. Huang72,58, K. X. Huang59, L. Q. Huang31,64, X. T. Huang50, Y. P. Huang1, Y. S. Huang59, T. Hussain74, F. Hölzken3, N. Hüsken35, N. in der Wiesche69, J. Jackson27, S. Janchiv32, J. H. Jeong10, Q. Ji1, Q. P. Ji19, W. Ji1,64, X. B. Ji1,64, X. L. Ji1,58, Y. Y. Ji50, X. Q. Jia50, Z. K. Jia72,58, D. Jiang1,64, H. B. Jiang77, P. C. Jiang46,h, S. S. Jiang39, T. J. Jiang16, X. S. Jiang1,58,64, Y. Jiang64, J. B. Jiao50, J. K. Jiao34, Z. Jiao23, S. Jin42, Y. Jin67, M. Q. Jing1,64, X. M. Jing64, T. Johansson76, S. Kabana33, N. Kalantar-Nayestanaki65, X. L. Kang9, X. S. Kang40, M. Kavatsyuk65, B. C. Ke81, V. Khachatryan27, A. Khoukaz69, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui1,64, N.  Kumar26, A. Kupsc44,76, W. Kühn37, J. J. Lane68, L. Lavezzi75A,75C, T. T. Lei72,58, Z. H. Lei72,58, M. Lellmann35, T. Lenz35, C. Li43, C. Li47, C. H. Li39, Cheng Li72,58, D. M. Li81, F. Li1,58, G. Li1, H. B. Li1,64, H. J. Li19, H. N. Li56,j, Hui Li43, J. R. Li61, J. S. Li59, K. Li1, K. L. Li19, L. J. Li1,64, L. K. Li1, Lei Li48, M. H. Li43, P. R. Li38,k,l, Q. M. Li1,64, Q. X. Li50, R. Li17,31, S. X. Li12, T.  Li50, W. D. Li1,64, W. G. Li1,a, X. Li1,64, X. H. Li72,58, X. L. Li50, X. Y. Li1,64, X. Z. Li59, Y. G. Li46,h, Z. J. Li59, Z. Y. Li79, C. Liang42, H. Liang1,64, H. Liang72,58, Y. F. Liang54, Y. T. Liang31,64, G. R. Liao14, Y. P. Liao1,64, J. Libby26, A.  Limphirat60, C. C. Lin55, D. X. Lin31,64, T. Lin1, B. J. Liu1, B. X. Liu77, C. Liu34, C. X. Liu1, F. Liu1, F. H. Liu53, Feng Liu6, G. M. Liu56,j, H. Liu38,k,l, H. B. Liu15, H. H. Liu1, H. M. Liu1,64, Huihui Liu21, J. B. Liu72,58, J. Y. Liu1,64, K. Liu38,k,l, K. Y. Liu40, Ke Liu22, L. Liu72,58, L. C. Liu43, Lu Liu43, M. H. Liu12,g, P. L. Liu1, Q. Liu64, S. B. Liu72,58, T. Liu12,g, W. K. Liu43, W. M. Liu72,58, X. Liu38,k,l, X. Liu39, Y. Liu81, Y. Liu38,k,l, Y. B. Liu43, Z. A. Liu1,58,64, Z. D. Liu9, Z. Q. Liu50, X. C. Lou1,58,64, F. X. Lu59, H. J. Lu23, J. G. Lu1,58, X. L. Lu1, Y. Lu7, Y. P. Lu1,58, Z. H. Lu1,64, C. L. Luo41, J. R. Luo59, M. X. Luo80, T. Luo12,g, X. L. Luo1,58, X. R. Lyu64, Y. F. Lyu43, F. C. Ma40, H. Ma79, H. L. Ma1, J. L. Ma1,64, L. L. Ma50, L. R. Ma67, M. M. Ma1,64, Q. M. Ma1, R. Q. Ma1,64, T. Ma72,58, X. T. Ma1,64, X. Y. Ma1,58, Y. M. Ma31, F. E. Maas18, I. MacKay70, M. Maggiora75A,75C, S. Malde70, Y. J. Mao46,h, Z. P. Mao1, S. Marcello75A,75C, Z. X. Meng67, J. G. Messchendorp13,65, G. Mezzadri29A, H. Miao1,64, T. J. Min42, R. E. Mitchell27, X. H. Mo1,58,64, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie20, I. B. Nikolaev4,c, Z. Ning1,58, S. Nisar11,m, Q. L. Niu38,k,l, W. D. Niu55, Y. Niu 50, S. L. Olsen64, S. L. Olsen10,64, Q. Ouyang1,58,64, S. Pacetti28B,28C, X. Pan55, Y. Pan57, A.  Pathak34, Y. P. Pei72,58, M. Pelizaeus3, H. P. Peng72,58, Y. Y. Peng38,k,l, K. Peters13,e, J. L. Ping41, R. G. Ping1,64, S. Plura35, V. Prasad33, F. Z. Qi1, H. Qi72,58, H. R. Qi61, M. Qi42, T. Y. Qi12,g, S. Qian1,58, W. B. Qian64, C. F. Qiao64, X. K. Qiao81, J. J. Qin73, L. Q. Qin14, L. Y. Qin72,58, X. P. Qin12,g, X. S. Qin50, Z. H. Qin1,58, J. F. Qiu1, Z. H. Qu73, C. F. Redmer35, K. J. Ren39, A. Rivetti75C, M. Rolo75C, G. Rong1,64, Ch. Rosner18, M. Q. Ruan1,58, S. N. Ruan43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning76, M. Scodeggio29A, K. Y. Shan12,g, W. Shan24, X. Y. Shan72,58, Z. J. Shang38,k,l, J. F. Shangguan16, L. G. Shao1,64, M. Shao72,58, C. P. Shen12,g, H. F. Shen1,8, W. H. Shen64, X. Y. Shen1,64, B. A. Shi64, H. Shi72,58, H. C. Shi72,58, J. L. Shi12,g, J. Y. Shi1, Q. Q. Shi55, S. Y. Shi73, X. Shi1,58, J. J. Song19, T. Z. Song59, W. M. Song34,1, Y.  J. Song12,g, Y. X. Song46,h,n, S. Sosio75A,75C, S. Spataro75A,75C, F. Stieler35, S. S Su40, Y. J. Su64, G. B. Sun77, G. X. Sun1, H. Sun64, H. K. Sun1, J. F. Sun19, K. Sun61, L. Sun77, S. S. Sun1,64, T. Sun51,f, W. Y. Sun34, Y. Sun9, Y. J. Sun72,58, Y. Z. Sun1, Z. Q. Sun1,64, Z. T. Sun50, C. J. Tang54, G. Y. Tang1, J. Tang59, M. Tang72,58, Y. A. Tang77, L. Y. Tao73, Q. T. Tao25,i, M. Tat70, J. X. Teng72,58, V. Thoren76, W. H. Tian59, Y. Tian31,64, Z. F. Tian77, I. Uman62B, Y. Wan55, S. J. Wang 50, B. Wang1, B. L. Wang64, Bo Wang72,58, D. Y. Wang46,h, F. Wang73, H. J. Wang38,k,l, J. J. Wang77, J. P. Wang 50, K. Wang1,58, L. L. Wang1, M. Wang50, N. Y. Wang64, S. Wang38,k,l, S. Wang12,g, T.  Wang12,g, T. J. Wang43, W.  Wang73, W. Wang59, W. P. Wang35,58,72,o, X. Wang46,h, X. F. Wang38,k,l, X. J. Wang39, X. L. Wang12,g, X. N. Wang1, Y. Wang61, Y. D. Wang45, Y. F. Wang1,58,64, Y. L. Wang19, Y. N. Wang45, Y. Q. Wang1, Yaqian Wang17, Yi Wang61, Z. Wang1,58, Z. L.  Wang73, Z. Y. Wang1,64, Ziyi Wang64, D. H. Wei14, F. Weidner69, S. P. Wen1, Y. R. Wen39, U. Wiedner3, G. Wilkinson70, M. Wolke76, L. Wollenberg3, C. Wu39, J. F. Wu1,8, L. H. Wu1, L. J. Wu1,64, X. Wu12,g, X. H. Wu34, Y. Wu72,58, Y. H. Wu55, Y. J. Wu31, Z. Wu1,58, L. Xia72,58, X. M. Xian39, B. H. Xiang1,64, T. Xiang46,h, D. Xiao38,k,l, G. Y. Xiao42, S. Y. Xiao1, Y.  L. Xiao12,g, Z. J. Xiao41, C. Xie42, X. H. Xie46,h, Y. Xie50, Y. G. Xie1,58, Y. H. Xie6, Z. P. Xie72,58, T. Y. Xing1,64, C. F. Xu1,64, C. J. Xu59, G. F. Xu1, H. Y. Xu67,2,p, M. Xu72,58, Q. J. Xu16, Q. N. Xu30, W. Xu1, W. L. Xu67, X. P. Xu55, Y. Xu40, Y. C. Xu78, Z. S. Xu64, F. Yan12,g, L. Yan12,g, W. B. Yan72,58, W. C. Yan81, X. Q. Yan1,64, H. J. Yang51,f, H. L. Yang34, H. X. Yang1, T. Yang1, Y. Yang12,g, Y. F. Yang1,64, Y. F. Yang43, Y. X. Yang1,64, Z. W. Yang38,k,l, Z. P. Yao50, M. Ye1,58, M. H. Ye8, J. H. Yin1, Junhao Yin43, Z. Y. You59, B. X. Yu1,58,64, C. X. Yu43, G. Yu1,64, J. S. Yu25,i, M. C. Yu40, T. Yu73, X. D. Yu46,h, Y. C. Yu81, C. Z. Yuan1,64, J. Yuan34, J. Yuan45, L. Yuan2, S. C. Yuan1,64, Y. Yuan1,64, Z. Y. Yuan59, C. X. Yue39, A. A. Zafar74, F. R. Zeng50, S. H. Zeng63A,63B,63C,63D, X. Zeng12,g, Y. Zeng25,i, Y. J. Zeng59, Y. J. Zeng1,64, X. Y. Zhai34, Y. C. Zhai50, Y. H. Zhan59, A. Q. Zhang1,64, B. L. Zhang1,64, B. X. Zhang1, D. H. Zhang43, G. Y. Zhang19, H. Zhang81, H. Zhang72,58, H. C. Zhang1,58,64, H. H. Zhang34, H. H. Zhang59, H. Q. Zhang1,58,64, H. R. Zhang72,58, H. Y. Zhang1,58, J. Zhang59, J. Zhang81, J. J. Zhang52, J. L. Zhang20, J. Q. Zhang41, J. S. Zhang12,g, J. W. Zhang1,58,64, J. X. Zhang38,k,l, J. Y. Zhang1, J. Z. Zhang1,64, Jianyu Zhang64, L. M. Zhang61, Lei Zhang42, P. Zhang1,64, Q. Y. Zhang34, R. Y. Zhang38,k,l, S. H. Zhang1,64, Shulei Zhang25,i, X. M. Zhang1, X. Y Zhang40, X. Y. Zhang50, Y. Zhang1, Y.  Zhang73, Y.  T. Zhang81, Y. H. Zhang1,58, Y. M. Zhang39, Yan Zhang72,58, Z. D. Zhang1, Z. H. Zhang1, Z. L. Zhang34, Z. Y. Zhang77, Z. Y. Zhang43, Z. Z.  Zhang45, G. Zhao1, J. Y. Zhao1,64, J. Z. Zhao1,58, L. Zhao1, Lei Zhao72,58, M. G. Zhao43, N. Zhao79, R. P. Zhao64, S. J. Zhao81, Y. B. Zhao1,58, Y. X. Zhao31,64, Z. G. Zhao72,58, A. Zhemchugov36,b, B. Zheng73, B. M. Zheng34, J. P. Zheng1,58, W. J. Zheng1,64, Y. H. Zheng64, B. Zhong41, X. Zhong59, H.  Zhou50, J. Y. Zhou34, L. P. Zhou1,64, S.  Zhou6, X. Zhou77, X. K. Zhou6, X. R. Zhou72,58, X. Y. Zhou39, Y. Z. Zhou12,g, Z. C. Zhou20, A. N. Zhu64, J. Zhu43, K. Zhu1, K. J. Zhu1,58,64, K. S. Zhu12,g, L. Zhu34, L. X. Zhu64, S. H. Zhu71, T. J. Zhu12,g, W. D. Zhu41, Y. C. Zhu72,58, Z. A. Zhu1,64, J. H. Zou1, J. Zu72,58
(BESIII Collaboration)
1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Bristol, H H Wills Physics Laboratory, Tyndall Avenue, Bristol, BS8 1TL, UK
64 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
65 University of Groningen, NL-9747 AA Groningen, The Netherlands
66 University of Hawaii, Honolulu, Hawaii 96822, USA
67 University of Jinan, Jinan 250022, People’s Republic of China
68 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
69 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
70 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
71 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
72 University of Science and Technology of China, Hefei 230026, People’s Republic of China
73 University of South China, Hengyang 421001, People’s Republic of China
74 University of the Punjab, Lahore-54590, Pakistan
75 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
76 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
77 Wuhan University, Wuhan 430072, People’s Republic of China
78 Yantai University, Yantai 264005, People’s Republic of China
79 Yunnan University, Kunming 650500, People’s Republic of China
80 Zhejiang University, Hangzhou 310027, People’s Republic of China
81 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at School of Physics, Beihang University, Beijing 100191 , China
Abstract
By analyzing 7.93 $\rm fb^{-1}$ of $e^{+}e^{-}$ collision data collected at the center-of-mass energy of 3.773 GeV with the BESIII detector operated at the BEPCII collider, we report the observation of the semimuonic decays of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$ with statistical significances of $12.5\sigma$ and $6.0\sigma$, respectively. Their decay branching fractions are determined to be ${\mathcal{B}}[D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}]=(2.36\pm 0.20^{+0.18}_{-0.27}\pm 0.48)\times 10^{-3}$ and ${\mathcal{B}}[D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}]=(0.78\pm 0.11^{+0.05}_{-0.09}\pm 0.15)\times 10^{-3}$, where the first and second uncertainties are statistical and systematic, respectively, and the third originates from the input branching fraction of $\bar{K}_{1}(1270)^{0}\to K^{-}\pi^{+}\pi^{0}$ or $K_{1}(1270)^{-}\to K^{-}\pi^{+}\pi^{-}$. Combining our branching fractions with the previous measurements of ${\mathcal{B}}[D^{+}\to\bar{K}_{1}(1270)^{0}e^{+}\nu_{e}]$ and ${\mathcal{B}}[D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}]$, we determine the branching fraction ratios to be ${\mathcal{B}}[D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}]/{\mathcal{B}}[D^{+}\to\bar{K}_{1}(1270)^{0}e^{+}\nu_{e}]=1.03\pm 0.14\begin{subarray}{c}+0.11\\ -0.15\end{subarray}$ and ${\mathcal{B}}[D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}]/{\mathcal{B}}[D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}]=0.74\pm 0.13\begin{subarray}{c}+0.08\\ -0.13\end{subarray}$. Using the branching fractions measured in this work and the world-average lifetimes of the $D^{+}$ and $D^{0}$ mesons, we determine the semimuonic partial decay width ratio to be $\Gamma[D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}]/\Gamma[D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}]=1.22\pm 0.10\begin{subarray}{c}+0.06\\ -0.09\end{subarray}$, which is consistent with unity as predicted by isospin conservation.
Experimental studies of the semileptonic decays of charmed mesons are important to deeply understand nonperturbative strong-interaction dynamics in weak decays. The semileptonic $D$ transitions into $\bar{K}_{1}(1270)$ offer an clean window to access the $K_{1}(1270)$ and $K_{1}(1400)$, which are mixtures of the $1^{3}P_{1}$ and $1^{1}P_{1}$ states of $K_{1}$ with mixing angle $\theta_{K_{1}}$. The determination of $\theta_{K_{1}}$ is essential for the theoretical description of $\tau$ Suzuki:1993yc , $B$ Hatanaka:2008xj ; Cheng:2004yj , and $D$ Cheng:2010vk ; Cheng:2003bn decays into axial-vector strange mesons. Throughout this Letter, charged-conjugate modes are implied and $K_{1}$ denotes $K_{1}(1270)$ unless stated otherwise.
Branching fractions (BF) for $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}(\ell=e$ or $\mu)$ were calculated in the Isgur-Scora-Grinstein-Wise (ISGW) quark model Isgur:1988gb and its update with ISGW2 Scora:1995ty . The BFs for $D^{+}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ and $D^{0}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ were predicted to be 0.1% and 0.3%, respectively. However, this model ignores mixing between ${}^{1}\rm P_{1}$ and ${}^{3}\rm P_{1}$ states. Recently, the BFs of $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ were calculated with other theoretical approaches, including three-point QCD sum rules (3PSR) Khosravi:2008jw , the covariant light-front quark model (CLFQM) Cheng:2017pcq , and light-cone QCD sum rules (LCSR) Momeni:2019uag . In general, the predicted BFs range from $10^{-3}$ to $10^{-2}$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf , and are sensitive to both the amplitude and sign of $\theta_{K_{1}}$. Measurements of the BFs of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ are important to test these theoretical calculations, to explore the nature of axial-vector strange mesons, and to understand the weak-decay mechanisms of $D$ mesons.
Lepton flavor universality (LFU) is a basic feature of the standard model (SM), in which the couplings between the three families of leptons and the gauge bosons are independent of lepton family. Nevertheless, a few hints of tension between experimental measurements and SM predictions have been reported in some cases, including semileptonic (SL) $B$ decays HFLAV and the Cabibbo-angle anomaly Coutinho:2019aiy ; Crivellin:2020lzu . Intensive tests of LFU in different SL decays of heavy mesons are important to understand these anomalies. Reference Fajfer2015 notes that there may indeed be observable LFU violation effects in the SL decays mediated via $c\to s\ell^{+}\nu_{\ell}$. In the SM, the ratio ${\mathcal{R}}^{\bar{K}_{1}}_{\mu/e}={\mathcal{B}}_{D\to\bar{K}_{1}\mu^{+}\nu_{\mu}}/{\mathcal{B}}_{D\to\bar{K}_{1}e^{+}\nu_{e}}$ is predicted to be $0.95-0.99$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf . To date, no experimental study of the semimuonic $D^{+(0)}$ decays into axial-vector mesons has been reported, and possible LFU in the SL $D^{+(0)}$ decays into axial-vector mesons has never been tested Ke:2023qzc . Observation of the decays and measurements of the BFs of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ offer an important opportunity to test $\mu$-$e$ LFU.
This paper reports the first observation of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ by using an $e^{+}e^{-}$ collision data sample corresponding to an integrated luminosity of 7.93 fb-1 BESIII:2024luminosity , recorded at the center-of-mass energy of $\sqrt{s}=3.773$ $\rm\,GeV$ with the BESIII detector BESIII . Using our BFs and the world average BFs of $D^{+}\to\bar{K}_{1}^{0}e^{+}\nu_{e}$, we provide the first test of the $\mu$-$e$ LFU with $D^{+}\to\bar{K}_{1}^{0}\ell^{+}\nu_{\ell}$. Combining the BFs of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ with the lifetimes of the $D^{+}$ and $D^{0}$, we test isospin invariance in $D\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ decays.
A description of the design and performance of the BESIII detector can be found in Ref. BESIII . The end cap TOF system was upgraded in 2015 using multi-gap resistive plate chamber technology, providing a time resolution of 60 ps BESIII:TOF , which benefits 67% of the data used in this analysis.
Simulated data samples are produced with a geant4-based geant4 Monte Carlo (MC) toolkit including the geometric description of the BESIII detector and the detector response. The simulation includes the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc kkmc . In the MC simulation, the production of open-charm processes directly produced via $e^{+}e^{-}$ annihilations are modelled with the generator conexc conexc . The ISR production of vector charmonium(-like) states and the continuum processes are incorporated in kkmc kkmc . All particle decays are modelled with evtgen evtgen using BFs either taken from the Particle Data Group pdg2022 , when available, or otherwise estimated with lundcharm lundcharm . Final state radiation from charged final state particles is incorporated using photos photos . The signal MC events of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}\mu^{+}\nu_{\mu}$ are generated with a special MC generator developed from the amplitude analysis of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}e^{+}\nu_{e}$ BESIII:k1ff .
This work analyzes the $e^{+}e^{-}\to\psi(3770)\to D\bar{D}$ decay chain. The single-tag (ST) candidate events are selected by reconstructing a $D^{-}$ or $\bar{D}^{0}$ in the hadronic final states $D^{-}\to K^{+}\pi^{-}\pi^{-}$, $K^{0}_{S}\pi^{-}$, $K^{+}\pi^{-}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{-}\pi^{0}$, $K^{0}_{S}\pi^{+}\pi^{-}\pi^{-}$, $K^{+}K^{-}\pi^{-}$, or $\bar{D}^{0}\to K^{+}\pi^{-}$, $K^{+}\pi^{-}\pi^{0}$, $K^{+}\pi^{-}\pi^{-}\pi^{+}$. These inclusively selected candidates are referred to as ST $D^{-}$ ($\bar{D}^{0}$) mesons. In the presence of the ST $D^{-}$ ($\bar{D}^{0}$) mesons, candidates for the signal decays are selected to form double-tag (DT) events. The BF of the semimuonic decay is determined by
${\mathcal{B}}_{\rm SL}=N_{\mathrm{DT}}/(N_{\mathrm{ST}}^{\rm tot}\,\cdot\varepsilon_{\rm sig}\cdot\mathcal{B}_{K_{1}}),$ (1)
where $N_{\rm ST}^{\rm tot}$ and $N_{\rm DT}$ are the total ST and DT yields in the data sample, $\varepsilon_{\rm sig}=\Sigma_{i}[(\varepsilon^{i}_{\rm DT}\,N^{i}_{\rm ST})/(\varepsilon^{i}_{\rm ST}\,N^{\rm tot}_{\rm ST})]$ is the efficiency of detecting the semimuonic decay in the presence of the ST $D^{-}$ ($\bar{D}^{0}$) meson and $\mathcal{B}_{K_{1}}$ is the BF of $\bar{K}_{1}\to K^{-}\pi^{+}\pi^{0(-)}$. Here, $i$ denotes the ST mode, and $\varepsilon_{\rm ST}$ and $\varepsilon_{\rm DT}$ are efficiencies for selecting the ST and DT candidates, respectively.
The selection criteria of $K^{\pm}$, $\pi^{\pm}$, $K^{0}_{S}$, and $\pi^{0}$ are the same as those used in Ref. BESIII:st . The ST $\bar{D}$ ($\bar{D}$ denotes $D^{-}$ or $\bar{D}^{0}$) mesons are selected based on two kinematic variables in the $e^{+}e^{-}$ center-of-mass frame: the energy difference $\Delta E\equiv E_{\bar{D}}-E_{\mathrm{beam}}$ and the beam-constrained mass $M_{\rm BC}\equiv\sqrt{E_{\mathrm{beam}}^{2}/c^{4}-|\vec{p}_{\bar{D}}|^{2}/c^{2}}$. Here, $E_{\mathrm{beam}}$ is the beam energy, and $E_{\bar{D}}$ and $\vec{p}_{\bar{D}}$ are the total energy and momentum of the $\bar{D}$ candidate, respectively. If there are multiple combinations for a given ST mode, the one giving the minimum $|\Delta E|$ is retained.
For each tag mode, the yield of reconstructed $\bar{D}$ mesons is obtained from an unbinned maximum likelihood fit to its $M_{\rm BC}$ distribution. In the fit, the signal distribution is modeled by the simulated shape convolved with a double-Gaussian resolution function to model the resolution difference between the data and the MC simulation. The combinatorial backgrounds are parameterized by an ARGUS function ARGUS with the endpoint fixed at $E_{\rm beam}=$1.8865 GeV.
The candidates with $M_{\rm BC}$ within (1.863,1.877) GeV/$c^{2}$ for ST $D^{-}$ and (1.859,1.873) GeV/$c^{2}$ for ST $\bar{D}^{0}$ are kept for further analyses. For detailed information about the $\Delta E$ requirements, the ST yields in data, and the ST efficiencies, see Table 1 of the supplemental material supplement . Summing over all tag modes, we obtain the total yields of ST $D^{-}$ and $\bar{D}^{0}$ mesons to be $(4149.9\pm 2.3_{\rm stat})\times 10^{3}$ and $(6306.8\pm 2.8_{\rm stat})\times 10^{3}$, respectively.
Candidates for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ or $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are reconstructed from the remaining tracks and showers not used in the $D^{+}$ or $D^{0}$ ST selection. The $\bar{K}_{1}^{0}$ or $K_{1}^{-}$ meson is reconstructed using its dominant decay $\bar{K}_{1}^{0}\to K^{-}\pi^{+}\pi^{0}$ or $K_{1}^{-}\to K^{-}\pi^{+}\pi^{-}$. Candidates for $K^{-}$, $\pi^{\pm}$ and $\pi^{0}$ are selected with the same criteria as those used in the ST selection. To identify muon candidates, we use the information from ${\rm d}E/{\rm d}x$, the time-of-flight system (TOF), and the electromagnetic calorimeter (EMC), and combined confidence levels for the muon, pion, and kaon hypotheses ($CL_{\mu}$, $CL_{\pi}$ and $CL_{K}$) are calculated. Muon candidates are required to satisfy $CL_{\mu}>0.001$, $CL_{\mu}>CL_{K}$ and $CL_{\mu}>CL_{\pi}$. To reduce the background from hadrons, muon candidates are also required to have a deposited energy in the EMC within less than 0.28 GeV. DT candidates must not contain any additional charged tracks, $N^{\rm char}_{\rm extra}=0$; otherwise they are vetoed.
To suppress the backgrounds containing extra $\pi^{0}$ mesons, such as the hadronic decays $D^{+}\to K^{-}\pi^{+}\pi^{+}\pi^{0}\pi^{0}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\pi^{+}\pi^{0}$, we require that there are no additional combinations of two photons that satisfy the requirements for a $\pi^{0}$ meson in the event selection, $N_{\rm extra}^{\pi^{0}}=0$. The maximum energy of any extra photons ($E_{\text{extra~{}}\gamma}^{\rm max}$) that have not been used in reconstructing the signal final states is required to be less than 0.20 GeV. The opening angle between the missing momentum and the most energetic shower, $\theta_{\vec{p}_{\mathrm{miss}},\gamma}$, is required to satisfy $\cos\theta_{\vec{p}_{\mathrm{miss}},\gamma}<0.69\,(0.54)$ for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$   ($D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$).
The presence of the undetectable neutrino is inferred via the kinematic quantity $U_{\mathrm{miss}}\equiv E_{\mathrm{miss}}-|\vec{p}_{\mathrm{miss}}|c$, where $E_{\mathrm{miss}}$ and $\vec{p}_{\mathrm{miss}}$ are the missing energy and momentum of the semimuonic candidate, respectively, calculated by $E_{\mathrm{miss}}\equiv E_{\mathrm{beam}}-\Sigma_{j}E_{j}$ and $\vec{p}_{\mathrm{miss}}\equiv\vec{p}_{D}-\Sigma_{j}\vec{p}_{j}$ in the $e^{+}e^{-}$ center-of-mass frame. The index $j$ sums over the $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$ and $\mu^{+}$ of the signal candidate, and $E_{j}$ and $\vec{p}_{j}$ are the energy and momentum of the $j$th particle, respectively. In order to improve the $U_{\mathrm{miss}}$ resolution, the $D$ energy is constrained to the beam energy and $\vec{p}_{D}\equiv-\hat{p}_{\bar{D}}\sqrt{E_{\mathrm{bceam}}^{2}/c^{2}-m_{D}^{2}c^{2}}$, where $\hat{p}_{\bar{D}}$ is the unit vector in the momentum direction of the ST $\bar{D}$, and $m_{D}$ is the nominal $D$ mass pdg2022 .
In order to suppress backgrounds from the hadronic decays $D^{+(0)}\to K^{-}\pi^{+}\pi^{+}\pi^{0(-)}$, with misidentification of a pion as a muon, the invariant mass of the $K^{-}\pi^{+}\pi^{0(-)}\mu^{+}$ combination ($M_{K^{-}\pi^{+}\pi^{0(-)}\mu^{+}}$), is required to be less than 1.72 (1.65) GeV/$c^{2}$. These selection criteria have been optimized by analyzing the inclusive MC samples with the figure-of-merit defined by $S/\sqrt{S+B}$, where $S$ and $B$ denote the signal and background yields, respectively.
Figure 1 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0(-)}}$ vs. $U_{\rm miss}$ of the accepted candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ in data. Selected $K_{1}$ candidates must have the invariant mass of the $K^{-}\pi^{+}\pi^{0(-)}$ combination within the interval $(1.163,1.343)$ GeV/$c^{2}$, which corresponds to about ±$\pm 1\Gamma$ (decay width) around the $K_{1}$ nominal mass pdg2022 .
Fig. 1: The $M_{K^{-}\pi^{+}\pi^{0(-)}}$ vs. $U_{\rm miss}$ distributions of the candidates for (left) $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and (right) $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$. The red lines denote the accepted mass range, $(1.163,1.343)$ GeV/$c^{2}$.
Figure 2 shows the distributions of $U_{\rm miss}$ of the accepted candidate events in data. Clear signals around zero are observed. The DT yields are extracted from unbinned extended maximum-likelihood fits to the $U_{\rm miss}$ distributions of data. In the fits, the signal shapes are described by the MC-simulated shape extracted from the signal MC events and convolved with a Gaussian function due to the resolution difference between the data and MC simulation. The combinatorial background shapes are modeled by the MC-simulated shape obtained from the inclusive MC samples. The numbers of signal, the combinatorial background events, and the Gaussian function are free parameters in the fits.
For $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$, the peaking background events from $D^{+}\to K^{-}\pi^{+}\pi^{+}\pi^{0}$ (BKG II) concentrate around 0 GeV; for $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$, the peaking background events from $D^{0}\to K^{-}\pi^{-}\eta^{\prime}_{\pi^{+}\pi^{-}\gamma}$ (BKG II), $D^{0}\to K^{-}\pi^{+}K^{+}\pi^{-}$ (BKG III), and $D^{0}\to K^{-}\pi^{+}\pi^{-}\pi^{+}\pi^{0}$ (BKG IV) concentrate around 0 GeV, 0 GeV, and 0.05 GeV, respectively. All peaking background components are modeled by the MC-simulated shapes obtained from the corresponding MC samples. The backgrounds peaking near zero are convolved with the same Gaussian smearing function as the signal with their yields fixed based on MC simulations. The peaking around 0.05 GeV is convolved with a separate Gaussian function and its yield and parameters of Gaussian function are free.
[…]
Fig. 2: Fits to the $U_{\rm miss}$ distributions of the candidates for (left) $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and (right) $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$. The points with error bars are data. The blue solid curves are the total fits, the green solid curves denote the semimuonic signal shapes, and the red solid curves are from the combinatorial background shapes (BKG I). In the left plot, the magenta dashed curve is the background contribution from $D^{+}\to K^{-}\pi^{+}\pi^{+}\pi^{0}$ (BKG II). In the right plot, the violet and magenta dashed curves are the background contributions from $D^{0}\to K^{-}\pi^{-}\eta^{\prime}_{\pi^{+}\pi^{-}\gamma}$ (BKG II) and $D^{0}\to K^{-}\pi^{+}K^{+}\pi^{-}$ (BKG III), respectively. The orange dashed curves is the background contribution from $D^{0}\to K^{-}\pi^{+}\pi^{-}\pi^{+}\pi^{0}$ (BKG IV).
The DT efficiencies, $\varepsilon^{i}_{\rm DT}$, obtained from signal MC samples with the same analysis procedures as for data, are summarized in Table 1 of Ref. supplement . The averaged signal efficiencies in the presence of the ST $\bar{D}$ mesons are $(8.45\pm 0.05)\%$ and $(10.92\pm 0.07)\%$ for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$, respectively. These efficiencies include correction factors which are discussed later. The final signal efficiency, $\varepsilon_{\rm sig}$, includes the BF of $\pi^{0}\to\gamma\gamma$ pdg2022 . The reliability of the MC simulation was verified by the agreement of the distributions of the momenta and $\cos\theta$ of $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$, and $\mu^{+}$ with the data.
By inserting $N_{\rm DT}$, $\varepsilon_{\rm sig}$, and $N_{\rm ST}^{\rm tot}$ into Eq. (1), we determine the product of ${\mathcal{B}}_{\rm SL}$ and ${\mathcal{B}}_{K_{1}}$ pdg2022 to be
${\mathcal{B}}^{+}_{\rm SL}{\mathcal{B}}^{0}_{K_{1}}=(10.92\pm 0.91\begin{subarray}{c}+0.82\\ -1.26\end{subarray})\times 10^{-4},$
${\mathcal{B}}^{0}_{\rm SL}{\mathcal{B}}^{-}_{K_{1}}=(2.62\pm 0.38\begin{subarray}{c}+0.17\\ -0.29\end{subarray})\times 10^{-4},$
where the first and second uncertainties are statistical and systematic. All related numbers are summarized in Table 1.
Table 1: The DT yields, $N_{\rm DT}$, in data, their statistical significances, the signal efficiencies, $\bar{\epsilon}_{\rm SL}$, and the obtained BFs, ${\mathcal{B}}_{\rm SL}$, for the two signal modes. The first and second uncertainties are statistical and systematic, respectively. The third uncertainty is from knowledge of the external BF, ${\mathcal{B}}_{K_{1}}$.
Signal decay $N_{\rm DT}$ Significance ($\sigma$) $\bar{\varepsilon}_{\rm SL}$ (%) ${\mathcal{B}}_{K_{1}}\,{\mathcal{B}}_{\rm SL}$ (×$\times 10^{-4}$) ${\mathcal{B}}_{K_{1}}$ (%) ${\mathcal{B}}_{\rm SL}$ (×$\times 10^{-3}$)
$D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ $382.8\pm 31.9$ $12.5$ $8.45\pm 0.05$ $10.92\pm 0.91\begin{subarray}{c}+0.82\\ -1.26\end{subarray}$ $46.24\pm 9.33$ $2.36\pm 0.20\begin{subarray}{c}+0.18\\ -0.27\end{subarray}\pm 0.48$
$D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ $180.3\pm 25.9$ $6.0$ $10.92\pm 0.07$ $2.62\pm 0.38\begin{subarray}{c}+0.17\\ -0.29\end{subarray}$ $33.74\pm 6.54$ $0.78\pm 0.11\begin{subarray}{c}+0.05\\ -0.09\end{subarray}\pm 0.15$
The systematic uncertainties in the BF measurement, which are assigned relative to the measured BF, are discussed below. The DT method ensures that most uncertainties arising from the ST selection cancel. The uncertainty associated with the ST yield $N_{\rm ST}^{\rm tot}$, is assigned as 0.3% after varying the signal and background shapes in the $M_{\rm BC}$ fit.
The uncertainties associated with the efficiencies of $\mu^{+}$ tracking, $K^{-}$ tracking and PID, $\pi^{+}$ tracking and PID, and $\pi^{0}$ reconstruction are investigated using data and MC samples of $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$ events and DT $D\bar{D}$ hadronic events ($\bar{D}^{0}\to K^{+}\pi^{-}$, $\bar{D}^{0}\to K^{+}\pi^{-}\pi^{0}$, and $\bar{D}^{0}\to K^{+}\pi^{+}\pi^{-}\pi^{-}$ vs. $D^{0}\to K^{-}\pi^{+}$, $D^{0}\to K^{-}\pi^{+}\pi^{+}\pi^{-}$, and $D^{0}\to K^{-}\pi^{+}\pi^{0}$, and $D^{-}\to K^{+}\pi^{-}\pi^{-}$ vs. $D^{+}\to K^{-}\pi^{+}\pi^{+}$). We assign the uncertainties associated with the $\mu^{+}$ tracking, $K^{-}$ tracking (PID), $\pi^{+}$ tracking (PID) and $\pi^{0}$ reconstruction to be 1.0%, 0.5% (0.5%), 0.5% (0.5%) and 2.0%, respectively.
In the studies of $\mu^{+}$ PID efficiencies, the 2D (momentum and $\cos\theta$) PID efficiencies of data and MC simulation of $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$ events are re-weighted to match those of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ decays. The efficiency differences are $-(1.5\pm 0.4)\%$ and $-(2.3\pm 0.4)\%$ for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$, respectively. The MC efficiencies are then corrected by these differences and used to determine the BFs of the semimuonic decays. After corrections, we assign the uncertainties associated with the $\mu^{+}$ PID to be 0.4% and 0.4%, respectively. The systematic uncertainties due to the combined requirements on $E_{\text{extra~{}}\gamma}^{\rm max},N_{\rm extra}^{\pi^{0}},N^{\rm char}_{\rm extra}$ and, separately, that on $\cos\theta_{\vec{p}_{\mathrm{miss}},\gamma}$ are investigated using the control sample of $D^{+(0)}\to K_{S}^{0}\pi^{0(-)}\mu^{+}\nu_{\mu}$. The corresponding systematic uncertainties are assigned as 3.9% and 4.0% for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ respectively; while those for $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are 2.6% and 2.3%, respectively.
To estimate the uncertainties due to the $M_{K^{-}\pi^{+}\pi^{0(-)}\mu^{+}}$ requirements, we use DT samples with the signal modes replaced by $D^{+}\to K^{0}_{S}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{0}_{S}\pi^{-}\mu^{+}\nu_{\mu}$ and assign uncertainties of 4.2% and 2.3%, for the $D^{+}$ and $D^{0}$ modes, respectively.
To estimate the uncertainties due to the $M_{K^{-}\pi^{+}\pi^{0(-)}}$ requirements, we re-measure the BFs with the nominal $K_{1}$ mass window half width of 0.90 GeV/$c^{2}$ widened to 1.10 GeV/$c^{2}$. The differences between the average BFs and the baseline values are assigned as the systematic uncertainties, which yields 0.6% for both decay modes.
The uncertainties associated with the fits to the $U_{\rm miss}$ distributions are estimated to be $\begin{subarray}{c}+0.5\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+4.4\%\\ -10.0\%\end{subarray}$) which are dominated by the uncertainty from the background shape. These systematic uncertainties are studied by altering the default MC background shape in three ways. Firstly, alternative MC samples are used to determine the background shape, where the relative fraction of the $q\bar{q}$ background is varied within the uncertainties BESqqbar . Secondly, the BFs of the major $D\bar{D}$ background sources are varied by their uncertainties pdg2022 . Thirdly, the non-$K_{1}^{-}(\bar{K}_{1}^{0})$ sources of $K^{-}\pi^{+}\pi^{-}(K^{-}\pi^{+}\pi^{0})$ ($\begin{subarray}{c}+0.0\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+0.0\%\\ -9.0\%\end{subarray}$)) are the dominant uncertainties. They are assigned to be the changes of the fitted DT yields after fixing non-resonant components by referring to the non-resonant fraction in $B\to J/\psi\bar{K}\pi\pi$ Belle:2002otd , since the non-resonant fraction in $D^{0(+)}\to\bar{K}\pi e^{+}\nu_{e}$ BESIII:2018jjm ; BESIII:2015hty is close to that in $B^{+}\to J/\psi\bar{K}\pi$ Belle:2010wrf .
The uncertainties related to the signal MC model are estimated to be 1.0% (1.0%), which correspond to the differences of the baseline signal efficiencies and those obtained with the signal MC samples after varying the input generator parameters by ±$\pm 1\sigma$. The uncertainty due to the limited MC statistics is 0.4% (0.4%).
The total systematic uncertainties are $\begin{subarray}{c}+7.5\%\\ -11.6\%\end{subarray}$ and $\begin{subarray}{c}+6.5\%\\ -11.1\%\end{subarray}$ for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$, respectively, obtained by adding all the individual contributions in quadrature. All systematic uncertainties are summarized in Table 2.
Table 2: Systematic uncertainties in % for the BF measurements.
Source $\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ $K_{1}^{-}\mu^{+}\nu_{\mu}$
$N^{\rm tot}_{\rm ST}$ 0.3 0.3
$\mu^{+}$ tracking 1.0 1.0
$K^{-},\pi^{\pm}$ tracking 1.0 1.5
$K^{-},\pi^{\pm}$ PID 1.0 1.5
$\pi^{0}$ reconstruction 2.0
$\mu^{+}$ PID 0.4 0.4
$E_{\rm extra\gamma}^{\rm max},N_{\pi^{0}}^{\rm extra},N^{\rm extra}_{\rm char}$ 3.9 2.6
$\cos\theta_{\vec{p}_{\mathrm{miss}},\gamma}$ 4.0 2.3
$M_{K^{-}\pi^{+}\pi^{0(-)}\mu^{+}}$ requirement 4.2 2.3
$M_{K^{-}\pi^{+}\pi^{0(-)}}$ requirement 0.6 0.6
$U_{\rm miss}$ fit $\begin{subarray}{c}+0.5\\ -8.8\end{subarray}$ $\begin{subarray}{c}+4.4\\ -10.0\end{subarray}$
Signal model 1.0 1.0
MC statistics 0.4 0.4
Total $\begin{subarray}{c}+7.5\\ -11.6\end{subarray}$ $\begin{subarray}{c}+6.5\\ -11.1\end{subarray}$
Using the world average of ${\mathcal{B}}_{K_{1}\to K^{-}\pi^{+}\pi^{0}}=(46.24\pm 9.33)\%$ K10 ${\mathcal{B}}_{K_{1}\to K^{-}\pi^{+}\pi^{-}}=(33.74\pm 6.54)\%$ K1m , we obtain:
$\small{\mathcal{B}}_{D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}}=(2.36\pm 0.20\begin{subarray}{c}+0.18\\ -0.27\end{subarray}\pm 0.48)\times 10^{-3},$
${\mathcal{B}}_{D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}}=(0.78\pm 0.11\begin{subarray}{c}+0.05\\ -0.09\end{subarray}\pm 0.15)\times 10^{-3},$
where the third uncertainty is from the input BFs.
To summarize, by analyzing 7.93 $\rm fb^{-1}$ of $e^{+}e^{-}$ collisions data taken at $\sqrt{s}=3.773$ GeV, we report the first observation of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ with statistical significances of $12.5\sigma$ and $6.0\sigma$, respectively. Their decay BFs are determined to be ${\mathcal{B}}[D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}]=(2.36\pm 0.20\begin{subarray}{c}+0.18\\ -0.27\end{subarray}\pm 0.48)\times 10^{-3}$ and ${\mathcal{B}}[D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}]=(0.78\pm 0.11\begin{subarray}{c}+0.05\\ -0.09\end{subarray}\pm 0.15)\times 10^{-3}$. Table 3 shows the comparison of the measured BFs with several theoretical calculations. Because the BFs predicted in theory are sensitive to the $K_{1}$ mixing angle, our BFs provides some information on the $K_{1}$ mixing angle.
Table 3: Comparison of the BFs measured in this work with the predicted values for $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ (×$\times 10^{-3}$) from various theories with differences given in units of standard deviations.
CLFQM Cheng:2017pcq 3PSR Khosravi:2008jw LCSR Momeni:2019uag LCSR Momeni:2019uag LCSR Momeni:2019uag CLFQM Bian:2021gwf LFQM Bian:2021gwf LCSR Bian:2021gwf This work
Decay $\theta_{K1}$ $33^{\circ}$ $37^{\circ}$ $-(34\pm 13)^{\circ}$ $45^{\circ}$
$D^{+}\to\bar{K}^{0}_{1}\mu^{+}\nu_{\mu}$ Predicted $2.60\pm 0.30$ $2.70\pm 0.25$ $18.59\pm 0.31$ $16.86\pm 0.27$ $19.73\pm 0.32$ $5.9\begin{subarray}{c}+0.47\\ -0.11\end{subarray}$ $3.7\begin{subarray}{c}+0.00\\ -0.41\end{subarray}$ $1.4\begin{subarray}{c}+0.15\\ -0.27\end{subarray}$ $2.36\pm 0.20\begin{subarray}{c}+0.18\\ -0.27\end{subarray}\pm 0.48$
Deviation ($\sigma$) 0.4 0.6 27.3 25.2 29.0 6.8 2.6 1.8
$D^{0}\to{K^{-}_{1}}\mu^{+}\nu_{\mu}$ Predicted $1.03\pm 0.1$ $8.09\pm 0.15$ $6.78\pm 0.12$ $8.92\pm 0.16$ $2.3\begin{subarray}{c}+0.19\\ -0.44\end{subarray}$ $1.5\begin{subarray}{c}+0.00\\ -0.16\end{subarray}$ $0.54\begin{subarray}{c}+0.06\\ -0.11\end{subarray}$ $0.78\pm 0.11\begin{subarray}{c}+0.05\\ -0.09\end{subarray}\pm 0.15$
Deviation ($\sigma$) 1.1 28.6 25.1 31.1 5.4 2.8 1.1
Combining the BFs measured in this work and the BFs of ${\mathcal{B}}[D^{+}\to\bar{K}_{1}^{0}e^{+}\nu_{e}]=(2.30\pm{0.26\begin{subarray}{c}+0.18\\ -0.21\end{subarray}}\pm 0.25)\times 10^{-3}$ BESIII:2019dptok1 and ${\mathcal{B}}[D^{0}\to K_{1}^{-}e^{+}\nu_{e}]=(1.09\pm{0.13\begin{subarray}{c}+0.09\\ -0.13\end{subarray}}\pm 0.12)\times 10^{-3}$ BESIII:2021dotok1 reported in the previous BESIII analysis, we determine the BF ratios ($\mathcal{R}^{D}_{\mu/e}=\frac{{\mathcal{B}}[D\to\bar{K}_{1}\mu^{+}\nu_{\mu}]}{{\mathcal{B}}[D\to\bar{K}_{1}e^{+}\nu_{e}]}$) to be $\mathcal{R}^{D^{+}}_{\mu/e}=1.03\pm 0.14\begin{subarray}{c}+0.11\\ -0.15\end{subarray}$ and $\mathcal{R}^{D^{0}}_{\mu/e}=0.74\pm 0.13\begin{subarray}{c}+0.08\\ -0.13\end{subarray}$. Here, the systematic uncertainties in ST yields, $K^{-}/\pi^{\pm}$ tracking and PID, and $\pi^{0}$ reconstruction cancel. These are consistent with $\mu$-$e$ lepton flavor universality in $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ decays within current uncertainties. Using the BFs of $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ and $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ measured in this work as well as the world-average lifetimes of the $D^{0}$ and $D^{+}$ mesons pdg2022 , we determine the partial decay width ratio $\Gamma[D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}]/\Gamma[D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}]=1.22\pm 0.10\begin{subarray}{c}+0.06\\ -0.09\end{subarray}$, which is consistent with unity as predicted by isospin conservation.
I Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2020YFA0406400, 2020YFA0406300; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.
Supplemental materials
Table 4 summarizes some analysis details for the single-tag $D^{-}$ candidates, namely the $\Delta E$ requirements, the ST yields in data, the ST efficiencies, and the double-tag efficiencies.
Table 4: The $\Delta E$ requirements, the obtained ST $\bar{D}^{0}(D^{-})$ yields ($N^{i}_{\rm ST}$) in data, the ST efficiencies ($\varepsilon^{i}_{\rm ST}$), and the DT efficiencies ($\varepsilon^{i}_{\rm DT}$). The efficiencies do not include the branching fractions of the $K^{0}_{S}$ and $\pi^{0}$ decays. The uncertainties are statistical only.
Tag mode $\Delta E$ (GeV) $N^{i}_{\rm ST}~{}(\times 10^{3})$ $\varepsilon^{i}_{\rm ST}~{}(\%)$ $\varepsilon^{i}_{\rm DT}~{}(\%)$
$D^{-}\to K^{+}\pi^{-}\pi^{-}$ $(-0.025,0.024)$ $2164.1\pm 1.6$ $51.17\pm 0.01$ $8.87\pm 0.09$
$D^{-}\to K_{S}^{0}\pi^{-}$ $(-0.025,0.026)$ $250.4\pm 0.5$ $50.74\pm 0.02$ $9.17\pm 0.09$
$D^{-}\to K^{+}\pi^{-}\pi^{-}\pi^{0}$ $(-0.057,0.046)$ $689.0\pm 1.2$ $25.50\pm 0.01$ $7.74\pm 0.12$
$D^{-}\to K_{S}^{0}\pi^{-}\pi^{0}$ $(-0.062,0.049)$ $558.5\pm 0.9$ $26.28\pm 0.01$ $8.47\pm 0.12$
$D^{-}\to K_{S}^{0}\pi^{-}\pi^{-}\pi^{+}$ $(-0.028,0.027)$ $300.5\pm 0.7$ $29.01\pm 0.01$ $8.13\pm 0.12$
$D^{-}\to K^{+}K^{-}\pi^{-}$ $(-0.024,0.023)$ $187.4\pm 0.5$ $41.06\pm 0.02$ $7.83\pm 0.10$
$\bar{D}^{0}\to K^{+}\pi^{-}$ $(-0.027,0.027)$ $1449.3\pm 1.3$ $65.34\pm 0.01$ $11.72\pm 0.09$
$\bar{D}^{0}\to K^{+}\pi^{-}\pi^{0}$ $(-0.062,0.049)$ $2913.2\pm 2.0$ $35.59\pm 0.01$ $11.16\pm 0.12$
$\bar{D}^{0}\to K^{+}\pi^{+}\pi^{-}\pi^{-}$ $(-0.026,0.024)$ $1944.2\pm 1.6$ $40.83\pm 0.01$ $9.95\pm 0.11$
Figure 3 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0}}$ from $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $M_{K^{-}\pi^{+}\pi^{-}}$ for $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ candidates, respectively. Results from the data and the inclusive MC sample are overlaid.
Fig. 3: Comparison of $M_{K^{-}\pi^{+}\pi^{0}}$ (left) and $M_{K^{-}\pi^{+}\pi^{-}}$ (right) of the candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$. The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples. A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.

### v2

@@
-Observation of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$
-Observation of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$
+Observation of $D\to\bar{K}_{1}(1270)\mu^{+}\nu_{\mu}$ and test of lepton flavor universality with $D\to\bar{K}_{1}(1270)\ell^{+}\nu_{\ell}$
+Observation of $D\to\bar{K}_{1}(1270)\mu^{+}\nu_{\mu}$ and test of lepton flavor universality with $D\to\bar{K}_{1}(1270)\ell^{+}\nu_{\ell}$
 M.
@@
 Using the branching fractions measured in this work and the world-average lifetimes of the $D^{+}$ and $D^{0}$ mesons, we determine the semimuonic partial decay width ratio to be $\Gamma[D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}]/\Gamma[D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}]=1.22\pm 0.10\begin{subarray}{c}+0.06\\ -0.09\end{subarray}$, which is consistent with unity as predicted by isospin conservation.
-Experimental studies of the semileptonic decays of charmed mesons are important to deeply understand nonperturbative strong-interaction dynamics in weak decays.
-The semileptonic $D$ transitions into $\bar{K}_{1}(1270)$ offer an clean window to access the $K_{1}(1270)$ and $K_{1}(1400)$, which are mixtures of the $1^{3}P_{1}$ and $1^{1}P_{1}$ states of $K_{1}$ with mixing angle $\theta_{K_{1}}$.
-The determination of $\theta_{K_{1}}$ is essential for the theoretical description of $\tau$ Suzuki:1993yc , $B$ Hatanaka:2008xj ; Cheng:2004yj , and $D$ Cheng:2010vk ; Cheng:2003bn decays into axial-vector strange mesons.
-Throughout this Letter, charged-conjugate modes are implied and $K_{1}$ denotes $K_{1}(1270)$ unless stated otherwise.
-Branching fractions (BF) for $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}(\ell=e$ or $\mu)$ were calculated in the Isgur-Scora-Grinstein-Wise (ISGW) quark model Isgur:1988gb and its update with ISGW2 Scora:1995ty .
+Experimental studies of the semileptonic (SL) decays of charmed mesons are important to deeply understand nonperturbative strong-interaction dynamics in weak decays.
+The SL $D$ transitions into $\bar{K}_{1}(1270)$ offer an clean window to access the $K_{1}(1270)$ and $K_{1}(1400)$, which are mixtures of the $1^{3}P_{1}$ and $1^{1}P_{1}$ states of $K_{1}$ with mixing angle $\theta_{K_{1}}$.
+The determination of $\theta_{K_{1}}$ is essential for the theoretical description of $\tau$ Suzuki:1993yc , $B$ Hatanaka:2008xj ; Cheng:2004yj , and $D$ Cheng:2010vk ; Wang:2020pyy ; Cheng:2003bn decays into axial-vector strange mesons.
+Throughout this paper, charged-conjugate modes are implied and $K_{1}$ denotes $K_{1}(1270)$ unless stated otherwise.
+Branching fractions (BFs) for $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}(\ell=e$ or $\mu)$ were calculated in the Isgur-Scora-Grinstein-Wise (ISGW) quark model Isgur:1988gb and its update with ISGW2 Scora:1995ty .
 The BFs for $D^{+}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ and $D^{0}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ were predicted to be 0.1% and 0.3%, respectively.
 However, this model ignores mixing between ${}^{1}\rm P_{1}$ and ${}^{3}\rm P_{1}$ states.
-Recently, the BFs of $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ were calculated with other theoretical approaches, including three-point QCD sum rules (3PSR) Khosravi:2008jw , the covariant light-front quark model (CLFQM) Cheng:2017pcq , and light-cone QCD sum rules (LCSR) Momeni:2019uag .
+Recently, the BFs of $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ were calculated with other theoretical approaches, including three-point quantum chromodynamics (QCD) sum rules (3PSR) Khosravi:2008jw , the covariant light-front quark model (CLFQM) Cheng:2017pcq , and light-cone QCD sum rules (LCSR) Momeni:2019uag .
 In general, the predicted BFs range from $10^{-3}$ to $10^{-2}$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf , and are sensitive to both the amplitude and sign of $\theta_{K_{1}}$.
@@
 Lepton flavor universality (LFU) is a basic feature of the standard model (SM), in which the couplings between the three families of leptons and the gauge bosons are independent of lepton family.
-Nevertheless, a few hints of tension between experimental measurements and SM predictions have been reported in some cases, including semileptonic (SL) $B$ decays HFLAV and the Cabibbo-angle anomaly Coutinho:2019aiy ; Crivellin:2020lzu .
+Nevertheless, a few hints of tension between experimental measurements and SM predictions have been reported in some cases, including SL $B$ decays HFLAV and the Cabibbo-angle anomaly Coutinho:2019aiy ; Crivellin:2020lzu .
 Intensive tests of LFU in different SL decays of heavy mesons are important to understand these anomalies.
@@
 In the SM, the ratio ${\mathcal{R}}^{\bar{K}_{1}}_{\mu/e}={\mathcal{B}}_{D\to\bar{K}_{1}\mu^{+}\nu_{\mu}}/{\mathcal{B}}_{D\to\bar{K}_{1}e^{+}\nu_{e}}$ is predicted to be $0.95-0.99$ Khosravi:2008jw ; Cheng:2017pcq ; Momeni:2019uag ; Bian:2021gwf .
-To date, no experimental study of the semimuonic $D^{+(0)}$ decays into axial-vector mesons has been reported, and possible LFU in the SL $D^{+(0)}$ decays into axial-vector mesons has never been tested Ke:2023qzc .
+To date, no experimental study of the semimuonic $D^{+(0)}$ decays into axial-vector mesons has been reported, and LFU in the SL $D^{+(0)}$ decays into axial-vector mesons has never been tested Ke:2023qzc .
 Observation of the decays and measurements of the BFs of $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ offer an important opportunity to test $\mu$-$e$ LFU.
@@
 BESIII .
-The end cap TOF system was upgraded in 2015 using multi-gap resistive plate chamber technology, providing a time resolution of 60 ps BESIII:TOF , which benefits 67% of the data used in this analysis.
+The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps BESIII:TOF , which benefits 67% of the data used in this analysis.
 Simulated data samples are produced with a geant4-based geant4 Monte Carlo (MC) toolkit including the geometric description of the BESIII detector and the detector response.
 The simulation includes the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc kkmc .
-In the MC simulation, the production of open-charm processes directly produced via $e^{+}e^{-}$ annihilations are modelled with the generator conexc conexc .
-The ISR production of vector charmonium(-like) states and the continuum processes are incorporated in kkmc kkmc .
-All particle decays are modelled with evtgen evtgen using BFs either taken from the Particle Data Group pdg2022 , when available, or otherwise estimated with lundcharm lundcharm .
+In the MC simulation, the production of open-charm processes directly produced via $e^{+}e^{-}$ annihilations are modeled with the generator conexc conexc .
+The ISR production of vector charmonium(like) states and the continuum processes are incorporated in kkmc kkmc .
+All particle decays are modeled with evtgen evtgen using BFs either taken from the Particle Data Group pdg2022 , when available, or otherwise estimated with lundcharm lundcharm .
 Final state radiation from charged final state particles is incorporated using photos photos .
@@
 In the fit, the signal distribution is modeled by the simulated shape convolved with a double-Gaussian resolution function to model the resolution difference between the data and the MC simulation.
-The combinatorial backgrounds are parameterized by an ARGUS function ARGUS with the endpoint fixed at $E_{\rm beam}=$1.8865 GeV.
+The combinatorial backgrounds are parametrized by an ARGUS function ARGUS with the endpoint fixed at $E_{\rm beam}=$1.8865 GeV.
 The candidates with $M_{\rm BC}$ within (1.863,1.877) GeV/$c^{2}$ for ST $D^{-}$ and (1.859,1.873) GeV/$c^{2}$ for ST $\bar{D}^{0}$ are kept for further analyses.
@@
 Summing over all tag modes, we obtain the total yields of ST $D^{-}$ and $\bar{D}^{0}$ mesons to be $(4149.9\pm 2.3_{\rm stat})\times 10^{3}$ and $(6306.8\pm 2.8_{\rm stat})\times 10^{3}$, respectively.
-Candidates for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ or $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are reconstructed from the remaining tracks and showers not used in the $D^{+}$ or $D^{0}$ ST selection.
+Candidates for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ or $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ are reconstructed from the remaining tracks and showers not used in the $D^{-}$ or $\bar{D}^{0}$ ST selection.
 The $\bar{K}_{1}^{0}$ or $K_{1}^{-}$ meson is reconstructed using its dominant decay $\bar{K}_{1}^{0}\to K^{-}\pi^{+}\pi^{0}$ or $K_{1}^{-}\to K^{-}\pi^{+}\pi^{-}$.
@@
 To reduce the background from hadrons, muon candidates are also required to have a deposited energy in the EMC within less than 0.28 GeV.
-DT candidates must not contain any additional charged tracks, $N^{\rm char}_{\rm extra}=0$; otherwise they are vetoed.
+The DT candidates must not contain any additional charged tracks, $N^{\rm char}_{\rm extra}=0$; otherwise they are vetoed.
 To suppress the backgrounds containing extra $\pi^{0}$ mesons, such as the hadronic decays $D^{+}\to K^{-}\pi^{+}\pi^{+}\pi^{0}\pi^{0}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\pi^{+}\pi^{0}$, we require that there are no additional combinations of two photons that satisfy the requirements for a $\pi^{0}$ meson in the event selection, $N_{\rm extra}^{\pi^{0}}=0$.
@@
 The index $j$ sums over the $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$ and $\mu^{+}$ of the signal candidate, and $E_{j}$ and $\vec{p}_{j}$ are the energy and momentum of the $j$th particle, respectively.
-In order to improve the $U_{\mathrm{miss}}$ resolution, the $D$ energy is constrained to the beam energy and $\vec{p}_{D}\equiv-\hat{p}_{\bar{D}}\sqrt{E_{\mathrm{bceam}}^{2}/c^{2}-m_{D}^{2}c^{2}}$, where $\hat{p}_{\bar{D}}$ is the unit vector in the momentum direction of the ST $\bar{D}$, and $m_{D}$ is the nominal $D$ mass pdg2022 .
+In order to improve the $U_{\mathrm{miss}}$ resolution, the $D$ energy is constrained to the beam energy and $\vec{p}_{D}\equiv-\hat{p}_{\bar{D}}\sqrt{E_{\mathrm{beam}}^{2}/c^{2}-m_{D}^{2}c^{2}}$, where $\hat{p}_{\bar{D}}$ is the unit vector in the momentum direction of the ST $\bar{D}$, and $m_{D}$ is the nominal $D$ mass pdg2022 .
 In order to suppress backgrounds from the hadronic decays $D^{+(0)}\to K^{-}\pi^{+}\pi^{+}\pi^{0(-)}$, with misidentification of a pion as a muon, the invariant mass of the $K^{-}\pi^{+}\pi^{0(-)}\mu^{+}$ combination ($M_{K^{-}\pi^{+}\pi^{0(-)}\mu^{+}}$), is required to be less than 1.72 (1.65) GeV/$c^{2}$.
@@
 $U_{\rm miss}$ of the accepted candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ in data.
-Selected $K_{1}$ candidates must have the invariant mass of the $K^{-}\pi^{+}\pi^{0(-)}$ combination within the interval $(1.163,1.343)$ GeV/$c^{2}$, which corresponds to about ±$\pm 1\Gamma$ (decay width) around the $K_{1}$ nominal mass pdg2022 .
+Figure 2 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0(-)}}$ of $D^{+(0)}\to K^{-}\pi^{+}\pi^{0(-)}\mu^{+}\nu_{\mu}$ candidates, in which significant $K_{1}(1270)$ signals can be seen.
+Selected $K_{1}$ candidates must have $M_{K^{-}\pi^{+}\pi^{0(-)}}$ within the interval $(1.163,1.343)$ GeV/$c^{2}$, which corresponds to about ±$\pm 1\Gamma$ (decay width) around the $K_{1}$ nominal mass pdg2022 .
 Fig.
@@
 The red lines denote the accepted mass range, $(1.163,1.343)$ GeV/$c^{2}$.
+Fig.
+2: The $M_{K^{-}\pi^{+}\pi^{0}}$ (left) and $M_{K^{-}\pi^{+}\pi^{-}}$ (right) distributions of the candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$.
+The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples.
+A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.
 Figure 3 shows the distributions of $U_{\rm miss}$ of the accepted candidate events in data.
@@
 The combinatorial background shapes are modeled by the MC-simulated shape obtained from the inclusive MC samples.
-The numbers of signal, the combinatorial background events, and the Gaussian function are free parameters in the fits.
+The signal yield, the combinatorial background yield, and the Gaussian function’s parameters are floated in the fits.
 For $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$, the peaking background events from $D^{+}\to K^{-}\pi^{+}\pi^{+}\pi^{0}$ (BKG II) concentrate around 0 GeV; for $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$, the peaking background events from $D^{0}\to K^{-}\pi^{-}\eta^{\prime}_{\pi^{+}\pi^{-}\gamma}$ (BKG II), $D^{0}\to K^{-}\pi^{+}K^{+}\pi^{-}$ (BKG III), and $D^{0}\to K^{-}\pi^{+}\pi^{-}\pi^{+}\pi^{0}$ (BKG IV) concentrate around 0 GeV, 0 GeV, and 0.05 GeV, respectively.
@@
 Fig.
-2: Fits to the $U_{\rm miss}$ distributions of the candidates for (left) $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and (right) $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$.
+3: Fits to the $U_{\rm miss}$ distributions of the candidates for (left) $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and (right) $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$.
 The points with error bars are data.
@@
 The final signal efficiency, $\varepsilon_{\rm sig}$, includes the BF of $\pi^{0}\to\gamma\gamma$ pdg2022 .
-The reliability of the MC simulation was verified by the agreement of the distributions of the momenta and $\cos\theta$ of $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$, and $\mu^{+}$ with the data.
+The reliability of the MC simulation has been verified by the agreement of the distributions of the momenta and $\cos\theta$ of $K^{-}$, $\pi^{+}$, $\pi^{0(-)}$, and $\mu^{+}$ with the data.
 By inserting $N_{\rm DT}$, $\varepsilon_{\rm sig}$, and $N_{\rm ST}^{\rm tot}$ into Eq. (1), we determine the product of ${\mathcal{B}}_{\rm SL}$ and ${\mathcal{B}}_{K_{1}}$ pdg2022 to be
@@
 ${\mathcal{B}}^{0}_{\rm SL}{\mathcal{B}}^{-}_{K_{1}}=(2.62\pm 0.38\begin{subarray}{c}+0.17\\ -0.29\end{subarray})\times 10^{-4},$
-where the first and second uncertainties are statistical and systematic. All related numbers are summarized in Table 1.
+where the first and second uncertainties are statistical and systematic, respectively. All related numbers are summarized in Table 1.
 Table 1: The DT yields, $N_{\rm DT}$, in data, their statistical significances, the signal efficiencies, $\bar{\epsilon}_{\rm SL}$, and the obtained BFs, ${\mathcal{B}}_{\rm SL}$, for the two signal modes.
@@
 $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ $180.3\pm 25.9$ $6.0$ $10.92\pm 0.07$ $2.62\pm 0.38\begin{subarray}{c}+0.17\\ -0.29\end{subarray}$ $33.74\pm 6.54$ $0.78\pm 0.11\begin{subarray}{c}+0.05\\ -0.09\end{subarray}\pm 0.15$
-The systematic uncertainties in the BF measurement, which are assigned relative to the measured BF, are discussed below.
+The systematic uncertainties in the BF measurement, which are assigned relative to the measured BFs, are discussed below.
 The DT method ensures that most uncertainties arising from the ST selection cancel.
@@
 We assign the uncertainties associated with the $\mu^{+}$ tracking, $K^{-}$ tracking (PID), $\pi^{+}$ tracking (PID) and $\pi^{0}$ reconstruction to be 1.0%, 0.5% (0.5%), 0.5% (0.5%) and 2.0%, respectively.
-In the studies of $\mu^{+}$ PID efficiencies, the 2D (momentum and $\cos\theta$) PID efficiencies of data and MC simulation of $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$ events are re-weighted to match those of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ decays.
+In the studies of $\mu^{+}$ PID efficiencies, the 2D (momentum and $\cos\theta$) PID efficiencies of data and MC simulation of $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$ events are reweighted to match those of $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ decays.
 The efficiency differences are $-(1.5\pm 0.4)\%$ and $-(2.3\pm 0.4)\%$ for $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$, respectively.
@@
 To estimate the uncertainties due to the $M_{K^{-}\pi^{+}\pi^{0(-)}\mu^{+}}$ requirements, we use DT samples with the signal modes replaced by $D^{+}\to K^{0}_{S}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{0}_{S}\pi^{-}\mu^{+}\nu_{\mu}$ and assign uncertainties of 4.2% and 2.3%, for the $D^{+}$ and $D^{0}$ modes, respectively.
-To estimate the uncertainties due to the $M_{K^{-}\pi^{+}\pi^{0(-)}}$ requirements, we re-measure the BFs with the nominal $K_{1}$ mass window half width of 0.90 GeV/$c^{2}$ widened to 1.10 GeV/$c^{2}$.
+To estimate the uncertainties due to the $M_{K^{-}\pi^{+}\pi^{0(-)}}$ requirements, we remeasure the BFs with the nominal $K_{1}$ mass window half width of 0.90 GeV/$c^{2}$ widened to 1.10 GeV/$c^{2}$.
 The differences between the average BFs and the baseline values are assigned as the systematic uncertainties, which yields 0.6% for both decay modes.
@@
 These systematic uncertainties are studied by altering the default MC background shape in three ways.
-Firstly, alternative MC samples are used to determine the background shape, where the relative fraction of the $q\bar{q}$ background is varied within the uncertainties BESqqbar .
-Secondly, the BFs of the major $D\bar{D}$ background sources are varied by their uncertainties pdg2022 .
-Thirdly, the non-$K_{1}^{-}(\bar{K}_{1}^{0})$ sources of $K^{-}\pi^{+}\pi^{-}(K^{-}\pi^{+}\pi^{0})$ ($\begin{subarray}{c}+0.0\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+0.0\%\\ -9.0\%\end{subarray}$)) are the dominant uncertainties.
-They are assigned to be the changes of the fitted DT yields after fixing non-resonant components by referring to the non-resonant fraction in $B\to J/\psi\bar{K}\pi\pi$ Belle:2002otd , since the non-resonant fraction in $D^{0(+)}\to\bar{K}\pi e^{+}\nu_{e}$ BESIII:2018jjm ; BESIII:2015hty is close to that in $B^{+}\to J/\psi\bar{K}\pi$ Belle:2010wrf .
+First, alternative MC samples are used to determine the background shape, where the relative fraction of the $q\bar{q}$ background is varied within the uncertainties BESqqbar .
+Second, the BFs of the major $D\bar{D}$ background sources are varied by their uncertainties pdg2022 .
+Third, the non-$K_{1}^{-}(\bar{K}_{1}^{0})$ sources of $K^{-}\pi^{+}\pi^{-}(K^{-}\pi^{+}\pi^{0})$ ($\begin{subarray}{c}+0.0\%\\ -8.8\%\end{subarray}$ ($\begin{subarray}{c}+0.0\%\\ -9.0\%\end{subarray}$)) are the dominant uncertainties.
+They are assigned to be the changes of the fitted DT yields after fixing nonresonant components by referring to the nonresonant fraction in $B\to J/\psi\bar{K}\pi\pi$ Belle:2010wrf , since the nonresonant fraction in $D^{0(+)}\to\bar{K}\pi e^{+}\nu_{e}$ BESIII:2018jjm ; BESIII:2015hty is close to that in $B^{+}\to J/\psi\bar{K}\pi$ Belle:2002otd .
 The uncertainties related to the signal MC model are estimated to be 1.0% (1.0%), which correspond to the differences of the baseline signal efficiencies and those obtained with the signal MC samples after varying the input generator parameters by ±$\pm 1\sigma$.
@@
 Because the BFs predicted in theory are sensitive to the $K_{1}$ mixing angle, our BFs provides some information on the $K_{1}$ mixing angle.
+Our BFs disfavor the theoretical calculations from LCSR Momeni:2019uag and CLFQM Bian:2021gwf , which are under the assumption of the mixing angle $\theta_{K_{1}}=-(34\pm 13)^{\circ}$, by more than $5\sigma$.
 Table 3: Comparison of the BFs measured in this work with the predicted values for $D^{+(0)}\to\bar{K}_{1}\mu^{+}\nu_{\mu}$ (×$\times 10^{-3}$) from various theories with differences given in units of standard deviations.
@@
 Here, the systematic uncertainties in ST yields, $K^{-}/\pi^{\pm}$ tracking and PID, and $\pi^{0}$ reconstruction cancel.
-These are consistent with $\mu$-$e$ lepton flavor universality in $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ decays within current uncertainties.
+These are consistent with theoretical predictions based on $\mu$-$e$ LFU in $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ decays within current uncertainties.
 Using the BFs of $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ and $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ measured in this work as well as the world-average lifetimes of the $D^{0}$ and $D^{+}$ mesons pdg2022 , we determine the partial decay width ratio $\Gamma[D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}]/\Gamma[D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}]=1.22\pm 0.10\begin{subarray}{c}+0.06\\ -0.09\end{subarray}$, which is consistent with unity as predicted by isospin conservation.
-I Acknowledgement
+I Acknowledgements
 The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support.
 This work is supported in part by National Key R&D Program of China under Contracts Nos.
-2023YFA1606000, 2020YFA0406400, 2020YFA0406300; National Natural Science Foundation of China (NSFC) under Contracts Nos.
+2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos.
 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No.
@@
 $\bar{D}^{0}\to K^{+}\pi^{+}\pi^{-}\pi^{-}$ $(-0.026,0.024)$ $1944.2\pm 1.6$ $40.83\pm 0.01$ $9.95\pm 0.11$
-Figure 3 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0}}$ from $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $M_{K^{-}\pi^{+}\pi^{-}}$ for $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ candidates, respectively.
+Figure 4 and 5 shows the distributions of $M_{K^{-}\pi^{+}}$, $M_{K^{-}\pi^{0(-)}}$ and $M_{\pi^{+}\pi^{0(-)}}$ for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ candidates, respectively.
 Results from the data and the inclusive MC sample are overlaid.
 Fig.
-3: Comparison of $M_{K^{-}\pi^{+}\pi^{0}}$ (left) and $M_{K^{-}\pi^{+}\pi^{-}}$ (right) of the candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$.
+4: Comparison of $M_{K^{-}\pi^{+}}$ (left), $M_{K^{-}\pi^{0}}$ (center) and $M_{\pi^{+}\pi^{0}}$ (right) of the candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$.
 The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples.
 A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.
+Fig.
+5: Comparison of $M_{K^{-}\pi^{+}}$ (left), $M_{K^{-}\pi^{-}}$ (center) and $M_{\pi^{+}\pi^{-}}$ (right) of the candidates for $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$.
+The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples.
+A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.